OpenStudy (anonymous):

1Amelia can make 4 shirts or 8 pound of food per week. Glenn can make 2 shirts or 8 pounds of food per week. 1)What is the opportunity cost of making a shirt for Amelia? For glenn? 2) what is the opportunity cost of making a poun of food for Amelia for Gleen? 3)who has an absoulte advantage in shirts? in food? 4)without trade, assume each person spends half their time making shirts and half making food. how mich of each good does each person supply and therefore consume without trade. 5) tell how could trade make each person better off. - what each person will produce

OpenStudy (anonymous):

1. Opportunity cost is found by how much will it cost me to buy 1 shirt compared to an amount of food? In this case, the opportunity cost of 4 shirts is 8 pounds of food But you need to find the opportunity cost of 1 shirt 1/4 = x/8 8/4 = x x =2 And therefore, your opportunity of one shirt = 2 pounds of food 2. The wording is confusing, but I think you're talking about comparative advantage. Check that up a little more. 3. Absolute advantage is the person who can make the most goods in each. Who can make the most shirts? ~ Amelia. Who can make the most food? ~ non of them have absolute advantage. 4. If it's half half, you can make what they're making now... so just add the two figures for Amelia and the two figures for Glenn 5. Not sure of the last one, although I think they should specialize in what they're best at.

OpenStudy (anonymous):

thank you. what specifically get for num #4??